Helicopter Rides.The price of each ticket is $150 and the variable costs average $60 per person. Steve's monthly fixed costs are $58,500. 1. What is Steve's contribution margin per ride? 2. How many tours must the company conduct in a month to break even? 3. What sales revenue is needed to produce a net profit of $36,000 per month? 4. What is Steve's contribution margin ratio? 5. If annual sales increase by $100,000, by what amount will operating income increase? 6. If annual sales decrease by $50,000, by what amount will operating income decrease?

I need help with #5 & #6
Answer #1

1) Contribution margin per ride = 150-60 = 90 per person

2) Break even = 58500/90 = 650

3) Required revenue = (58500+36000)/0.60 = $157500

4) Contribution margin ratio = 90/150 = 60%

5) Operating income increase by = 100000*60% = 60000

6) Operating income decrease by = 50000*60% = -30000